Solve the following equations.

a) x^3 + x = 4x^2 - 6

b) 3x^4 - 5x^2 + 2 = 0


1
Expert's answer
2021-11-23T14:03:30-0500

a) "x^3+x=4x^2\u22126"


Step 1: Subtract "4x^2-6" from both sides.

"x^3+x\u2212(4x^2\u22126)=4x^2\u22126\u2212(4x^2\u22126)"


"x^3\u22124x^2+x+6=0"


Step 2: Factor left side of equation.

"(x+1)(x\u22122)(x\u22123)=0"


Step 3: Set factors equal to 0.

"x+1=0"  or "x\u22122=0"  or "x\u22123=0"


"x=\u22121"  or "x=2" or "x=3"


b) "3x^4 - 5x^2 + 2 = 0"


Step 1: Factor left side of equation.


"(x+1)(x\u22121)(3x^2\u22122)=0"


Step 2: Set factors equal to 0.


"x+1=0" or "x\u22121=0"  or "3x^2\u22122" =0


"x=\u22121"  or "x=1" or "x=0.816496581"  or "x=\u22120.816496581"